Use the Euclidean algorithm to find gcd(2074, 2457) = d.


1
Expert's answer
2021-04-13T13:39:37-0400

Let us use the Euclidean algorithm to find "d=\gcd(2074, 2457)":


"2457=2074\cdot1+383\\
2074=383\cdot 5+159\\
383=159\cdot 2+65\\
159=65\cdot2+29\\
65=29\cdot 2+7\\
29=7\cdot4+1\\
7=7\cdot1+0"

We conclude that "d=\gcd(2074, 2457)=1."